OpenStudy (anonymous):

Simplify each expression. a. ln(1/e) b. ln(square root e) c. ln(e) d. ln(e^−9) e. (ln(e))^−9

myininaya (myininaya):

ln1-lne=0-1=-1 is a ln(e^(1/2))=1/2lne=1/2*1=1/2 is b lne=1 is c -9lne=-9*1=-9 is d (1)^(-9)=1 is e